OpenStudy (pphalke):

-8=7-(8-b) its saying write a solution set

OpenStudy (mathmale):

I see you've posted a number of questions here over the last hour. I urge you to think back about what you've learned and apply at least some of it to solving this new problem. What could you do to simplify -8=7-(8-b) ... so that you could more easily find the solution set?

OpenStudy (mathmale):

How would you move that '7' on the right side to the left side? Think: What'd be the proper arithmetic operation to accomplish that?

OpenStudy (pphalke):

i have no idea

OpenStudy (mathmale):

Method 1: -8=7-(8-b) Subtract 7 from both sides of this equation. Result: -15=-(8-b) Next, mult. the quantities inside the parentheses by -1: -15=-8+b Last, isolate b by adding 8 to both sides. Try it, please. Share your result.